# You Can Wash Hands Better: Accurate Daily Handwashing Assessment with Smartwatches

### v2

@@ 5.1 Data Acquisition
5.2 User-Dependent Results
(1) Overall Results. We report overall results including accuracy, precision, recall, and F1 score in Table III. The accuracy is computed via Equation 3.
where $N_{p}$ represents the length of the test time-series of the $p$-th participant; $S_{p,i}$ and $S_{p,i}^{*}$ represent the ground-truth and the UWash output on the $i$-th sample of the $p$-th participant, respectively; $I(S_{p,i}^{*}==S_{p,i})$ outputs 1 if UWash recognizes correctly on $S_{p,i}$, otherwise 0. The table shows that UWash can achieve the sample-wise classification accuracy of 86.31% directly. With simple yet efficient post smoothing methods, i.e., multiple test voting and the mode filter (see Section 4.2), UWash can eventually achieve an accuracy of 92.27%.
TABLE III: Results on all participants. The results show that UWash performs well and can be enhanced with post-smoothing methods of multiple test voting (MTV) and the mode filter (TMF). Moreover, UWash is only with $\sim$2.5parameters of ResNet-18-1D, much more lightweight. TS (two-stream), SE (Squeeze-and-Excitation Module), PPM (Pyramid Pooling Module). $\uparrow$ means the larger the better. $\downarrow$ means the less the better.
@@ 5.2 User-Dependent Results
Uwash(Ours) 92.27% 51 sample-wise Yes
Figure 7: Confusion Matrix of Gesture Prediction. UWash works well for all gestures, especially for the $2$nd, $3$rd, $6$th, and $7$th gestures (zoom in for a better view).
(2) Performance on Gestures. We show the confusion matrix of UWash on 10 gestures (9 handwashing gestures + 1 background) in Fig. 7. Though the data of these 10 gestures are not quite balanced (29.2% of background), UWash works well for all gestures, especially for the $2$nd, $3$rd, $6$th, and $7$th gestures. Besides, we find two types of false recognition are dominant. The first happens between the background and gestures of 1 and 9. We think this is because the background contains diverse activities such as wetting hands with water, applying soap, and drying hands with a towel, which may largely increase the difficulty to be classified correctly with its post-activity (G1) or pre-activity (G9). The other false happens between two successive gestures. As the start/end time of every gesture is annotated manually, the discordance across different labeling workers, participants, and locations may cause false between successive gestures.
(3) Performance on Participants. For the $p$-th participant, we use Equation 4 to compute the accuracy.
$Accuracy_{p}=\frac{\sum_{i=1}^{N_{p}}I(S_{p,i}^{*}==S_{p,i})}{N_{p}}$ (4)
where symbols share the same meanings with Equation 3.
Fig. 9 shows that 33 of 51 participants achieve an accuracy of over 95%; only 6 of them have an accuracy of less than 85%; the average accuracy is 92.33% (orange bar). The results demonstrate that, for the seen participants, UWash achieves sample-wise handwashing gesture recognition effectively.
Figure 9: Results on participants. 33 of 51 participants are more than 95%; 6 of 51 are less than 85%
(4) Start/End Detection. Deserved to reiterate, UWash can automatically detect the handwashing start/end time, not requiring to work along with additional sensors [16, 12, 8] or to be awakened manually [12]. We use Equation 5 to compute the start time detection error.
$Error_{s}=\left|t_{s}^{*}-t_{s}\right|$ (5)
[…]
@@
(5) Scoring Results. UWash is the first work to score handwashing following the WHO guidelines. We use methods described in Section 4.3 to compute handwashing scores on the test set. Further, we compute the Mean and SD of scoring errors against ground truth. As shown in Table V, the mean and the SD are less than 5 points, indicating UWash can score handwashing well.
In Fig. 8, we visualize how UWash works on a sample of the 26th participant. UWash outputs sample-wise gesture classification, with which we can detect the start/end time, estimate the duration of gestures, score gestures with the estimated duration, and score the whole handwashing procedure. Thus UWash can help users to promote their handwashing practice with the estimated scores in daily life.
@@ 5.3 Cross-Domain Results
TABLE VI: Cross-domain results. $\uparrow$ means the higher the better. $\downarrow$ means the less the better. CP, CP-CL, and CP-CL-CT are for Cross-Participant, Cross-Participant-Cross-Location, and Cross-Participant-Cross-Location-Cross-Time.
@@
(1) Cross-Participant. We train UWash with data of 50 out of 51 participants and test the trained model with data of the remaining participant. We conduct this leave-one-participant-out process over 51 participants respectively to evaluate the cross-participant performance. As shown in Fig. 10, accuracy among some left-out participants is more discrete than those in the user-dependent setting, shown in Fig. 9. For example, the 13th and the 38th participants have relatively special personal handwashing styles, resulting in an accuracy of $<$60%. This indicates that they may not follow the WHO guidelines. In this case, UWash could remind users to improve their hand hygiene techniques. As expected, the mean accuracy in the cross-participant setting decreases to 83.34%, since the personalized gestures of these individuals are not included in the training set.
@@
The experimental results are shown results in Fig. 11 and Table VI. In this setting, the performances have similar characteristics to those in the cross-participant setting, e.g., the accuracy is more discrete; the 13th and the 38th participants have the lowest accuracy; the mean accuracy drops to 81.45%; the start/end time detections are achieved well.
@@
(3) Cross-Participant-Cross-Location-Cross-Time.
Nine months after the data collection described in Sec. 5.1, we further randomly recruit 10 passersby in a hospital to wash their hands 5 times, and follow the same process to obtain a new dataset. We apply the trained model in user-dependent results on the new dataset to evaluate the cross-time performance.
Figure 11: Cross-Participant-Cross-Location Results.
TABLE VII: Cross-Participant-Cross-Location-Cross-Time Results. UWash is stable and promising across time.
@@ 6 Conclusion
In this paper, we introduce UWash, a smartwatch-only handwashing assessment system, to raise people’s awareness of handwashing in daily use and adherence to the WHO handwashing guidelines. UWash takes the data of the IMU sensors of smartwatches, i.e., accelerometers and gyroscopes, as inputs, feeds the inputs into a dual-branch U-Nets, and outputs sample-wise gesture recognition results effectively. UWash can detect handwashing start/end time, estimate the duration of every handwashing gesture, and score gestures as well as the entire procedure following WHO guidelines. Experimental results over 51 participants show that UWash works well in the user-dependent, cross-participant, and cross-participant-cross-location settings. In addition, UWash still performs promising on a dataset collected 9 months later in a hospital. Moreover, UWash models are lightweight, only 496KB, with great potential to be deployed on edge devices in the future.

### v3

@@ 5.1 Data Acquisition
TABLE III: Results on all participants. The results show that UWash performs well and can be enhanced with post-smoothing methods of multiple test voting (MTV) and the mode filter (TMF). Moreover, UWash is much more lightweight with $\sim$2.5parameters of ResNet-18-1D. Two-stream (TS), Squeeze-and-Excitation Module (SE), Pyramid Pooling Module (PPM). $\uparrow$ means the larger the better. $\downarrow$ means the less the better.
@@
5.2 Evaluation Metrics
(1) Gesture recognition accuracy. The accuracy is computed via Equation 3.
where $N_{p}$ represents the length of the test time-series of the $p$-th participant; $S_{p,i}$ and $S_{p,i}^{*}$ represent the ground-truth and the UWash output on the $i$-th sample of the $p$-th participant, respectively; $I(S_{p,i}^{*}==S_{p,i})$ outputs 1 if UWash recognizes correctly on $S_{p,i}$, otherwise 0.
(2) Participant-wise accuracy. For the $p$-th participant, we use Equation 4 to compute the recognition accuracy of her/him.
$Accuracy_{p}=\frac{\sum_{i=1}^{N_{p}}I(S_{p,i}^{*}==S_{p,i})}{N_{p}}$ (4)
where symbols share the same meanings with Equation 3.
(3) We also apply precision, recall, and F1 to evaluate the gesture recognition performance.
(4) Start/End Detection Error. We use Equation 5 to compute the start time detection error.
$Error_{s}=\left|t_{s}^{*}-t_{s}\right|$ (5)
[…]
(5) Scoring Error. UWash is the first work to score handwashing following the WHO guidelines. We use methods described in Section 4.3 to compute handwashing scores on the test set. We then use Equation 6 to compute the scoring error.
$Error_{s}=\left|\mathrm{score}^{*}-\mathrm{score}\right|$ (6)
where $\mathrm{score}^{*}$ and $\mathrm{score}$ represent the estimation and the ground-truth of handwashing scores, respectively.
5.3 User-Dependent Results
(1) Overall Results. We report overall results including accuracy, precision, recall, and F1 score in Table III. The table shows that UWash can achieve the sample-wise classification accuracy of 86.31% directly. With simple yet efficient post smoothing methods, i.e., multiple test voting and the mode filter (described in Section 4.2), UWash can eventually achieve an accuracy of 92.27%.
@@ 5.3 User-Dependent Results
Uwash(Ours) 92.27% 51+10+10 sample-wise Yes
Figure 7: Confusion Matrix of Gesture Prediction. UWash works well for all gestures, especially for the $2$nd, $3$rd, $6$th, and $7$th gestures.
@@
(2) Performance on Gestures. We show the confusion matrix of UWash on 10 gestures (9 handwashing gestures + 1 background) in Fig. 7. Though the data of these 10 gestures are not quite balanced (29.2% of background), UWash works well for all gestures, especially for the $2$nd, $3$rd, $6$th, and $7$th gestures. Besides, we find two types of false recognition are dominant. The first happens between the background and gestures of 1 and 9. We think this is because the background contains diverse activities such as wetting hands with water, applying soap, and drying hands with a towel, which may largely increase the difficulty to be classified correctly with its post-activity (G1) or pre-activity (G9). The other false recognition happens between two successive gestures. As the start/end time of every gesture is annotated manually, the discordance across different labeling workers, participants, and locations may cause false between successive gestures.
(3) Performance on Participants. Fig. 8 demonstrates the participant-wise accuracy, which shows that 33 of 51 participants achieve an accuracy of over 95%; only 6 of them have an accuracy of less than 85%; the average accuracy is 92.33% (orange bar). The results demonstrate that, for the seen participants, UWash achieves sample-wise handwashing gesture recognition effectively.
Figure 8: Results on participants. 33 out of 51 participants have an accuracy greater than 95%, while 6 out of 51 are less than 85%
(4) Start/End Detection. Deserved to reiterate, UWash can automatically detect the handwashing start/end time, not requiring to work along with additional sensors [18, 12, 8] or to be awakened manually [12]. We compute the start/end detection errors over the test set and report the Mean and standard deviation (SD) in Table V. The table shows that Means and SDs are all within 1 second, indicating UWash can detect handwashing events correctly and stably.
(5) Scoring Results. We report the handwashing scoring results in Table V. As shown in Table V, the mean and the SD are less than 5 points, indicating UWash can score handwashing well.
(6) Visualization. In Fig. 9, we visualize how UWash works on a sample of the 26th participant. UWash outputs sample-wise gesture classification, with which we can detect the start/end time, estimate the duration of gestures, score gestures with the estimated duration, and score the whole handwashing procedure. Thus UWash can help users to promote their handwashing practice with the estimated scores in daily life.
@@ 5.4 Cross-Domain Results
Figure 10: Cross-Participant Accuracy Results by Individual Participant: Accuracy for Different Participants
TABLE VI: Cross-domain results(Average Accuracy and Errors). $\uparrow$ means the higher the better. $\downarrow$ means the less the better. CP, CP-CL, and CP-CL-CT are for Cross-Participant, Cross-Participant-Cross-Location, and Cross-Participant-Cross-Location-Cross-Time.
@@
(1) Cross-Participant. We train UWash with data of 50 out of 51 participants and test the trained model with data of the remaining participant. We conduct this leave-one-participant-out process over 51 participants respectively to evaluate the cross-participant performance. As shown in Fig. 10, accuracy among some left-out participants is more discrete than those in the user-dependent setting. For example, the 13th and the 38th participants have relatively special personal handwashing styles, resulting in an accuracy of $<$60%. This indicates that they may not follow the WHO guidelines. In this case, UWash could remind users to improve their hand hygiene techniques. As expected, the mean accuracy in the cross-participant setting decreases to 83.34%, since the personalized gestures of these individuals are not included in the training set.
@@
The experimental results are shown in Fig. 11 and Table VI. In this setting, the performances have similar characteristics to those in the cross-participant setting, e.g., the accuracy is more discrete; the 13th and the 38th participants have the lowest accuracy; the mean accuracy drops to 81.45%; the start/end time detections are achieved well.
@@
(3) Cross-Participant-Cross-Location-Cross-Time. Nine months after the data collection described in Sec. 5.1, we further randomly recruit 10 passersby in a hospital to wash their hands 5 times, and follow the same process to obtain a new dataset. We apply the trained model in user-dependent results on the new dataset to evaluate the cross-time performance.
Figure 11: Cross-Participant-Cross-Location Accuracy by Individual Participant: Accuracy Distribution across Locations.
TABLE VII: Cross-Participant-Cross-Location-Cross-Time Results for Individual Participants. UWash is stable and promising across time.
@@
Figure 12: The Wild Study. UWash can work under the conditions when people shuffle the handwashing steps or omit steps.
(4) The Wild Study. To test UWash in the wild, we further recruit 10 passersby in a laboratory hall to evaluate their handwashing quality under three conditions: their usual handwashing method without prior training on the WHO-recommended technique (not trained), after receiving our training on the WHO handwashing guideline (trained), and when they intentionally shuffle the order of the WHO handwashing steps or omit certain steps (random). Five trials are collected for each individual and condition. We directly apply the UWash model trained in the in-domain evaluation section (Sec. 5.3) to estimate the handwashing scores, and the average scores are presented in Fig. 12.
As shown in Fig. 12, almost all not-trained participants have low handwashing scores, indicating that they do not adhere to the WHO-recommended handwashing procedure in their daily life. After training, their handwashing scores significantly improve, reaching an average score of approximately 85. When participants deliberately shuffle the handwashing steps or omit steps, UWash can detect a decline in their scores, demonstrating its effectiveness in identifying out-of-sequence guestures or missed steps. In practical use, UWash can alert users to any omitted steps, thereby enhancing handwashing quality in a targeted manner.
It is worth noting that participants P4 to P10 only shuffle the order of handwashing steps without omitting any, and their scores should have been close to those obtained during the trained session. However, these scores were slightly lower because, during the random session, they did not strictly adhere to the prescribed duration for handwashing. Since UWash evaluates scores based on the duration spent on each step, this deviation from the recommended durations leads to a reduction in overall scores.
@@ 6 Conclusion
In this paper, we introduce UWash, a smartwatch-only handwashing assessment system, to raise people’s awareness of handwashing in daily use and adherence to the WHO handwashing guidelines. UWash takes the data of the IMU sensors of smartwatches, i.e., accelerometers and gyroscopes, as inputs, feeds the inputs into a dual-branch U-Nets, and outputs sample-wise gesture recognition results effectively. UWash can detect handwashing start/end time, estimate the duration of every handwashing gesture, and score gestures as well as the entire procedure following WHO guidelines. Experimental results over 51 participants show that UWash works well in the user-dependent, cross-participant, and cross-participant-cross-location settings. In addition, UWash still performs promising 9 months later in a hospital and 2 years later in a hall during an in-wild test. Moreover, UWash models are lightweight, only 496KB, with great potential to be deployed on edge devices in the future.